Chapter 28, Problem 040 A wire 1.91 m long carries a current of 19.1 A and makes an angle of 41.8 with a uniform magnetic field of magnitude B = 1.96 T. Calculate the magnetic force on the wire. Number Units the tolerance is +/-7%
Answer #1

We know that

F = BILsin\theta

where \theta is the angle between current and field

F = 1.96*19.1*1.91*sin 41.8

F = 47.65 N